OpenStudy (anonymous):

how would i solve the equation log4(x-4)=-3? wouldnt i take 4 to the -3 exponent to start? im lost

OpenStudy (anonymous):

yes you would

OpenStudy (anonymous):

you would write \[4^{-3}=x-4\] and solve for x

OpenStudy (anonymous):

assuming of course that your problem is \[\log_4(x-4)=-3\]

OpenStudy (anonymous):

thanks for the response. so would that be x-4=.016?

OpenStudy (anonymous):

well lets see. \[4^{-3}=x-4\] \[x=4+4^{-3}=4+\frac{1}{64}\] whatever that is

OpenStudy (anonymous):

how would I find the domain of a function \[\log _{2\left( 25-x ^{2} \right)}\]? I think it is (5,-5)